OpenStudy (anonymous):

Perpendicular to y = 4x + 2 and passing through the point (0, -4)

OpenStudy (anonymous):

For perpendicular lines, you just flip the current slope your given and change the sign. You keep the same y-intercept. So you answer would be y=-1/4x+2

zepdrix (zepdrix):

Woops, while the slope part is true, we'll often have a `new` y-intercept! :) \[\Large\rm y=-\frac{1}{4}x+b\]Plug in the coordinate that they gave you to find your new y-intercept.\[\Large\rm -4=-\frac{1}{4}\cdot 0+b\]

zepdrix (zepdrix):

\(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

Hm. I was taught different, but I can see how you got that.